$SPAD/src/input hexadec.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool hexadec.output
)set message test on
)set message auto off
)clear all
--S 1 of 7
r := hex(22/7)
--R
--R
--R          ___
--R   (1)  3.249
--R                                                   Type: HexadecimalExpansion
--E 1
--S 2 of 7
r + hex(6/7)
--R
--R
--R   (2)  4
--R                                                   Type: HexadecimalExpansion
--E 2
--S 3 of 7
[hex(1/i) for i in 350..354]
--R
--R
--R   (3)
--R       ___________________________________________________
--R   [0.00BB3EE721A54D88, 0.00BAB6561, 0.00BA2E8, 0.00B9A7862A0FF465879D5F,
--R       _____________________________
--R    0.00B92143FA36F5E02E4850FE8DBD78]
--R                                             Type: List(HexadecimalExpansion)
--E 3
--S 4 of 7
hex(1/1007)
--R
--R
--R   (4)

--R   0.
--R     OVERBAR
--R        0041149783F0BF2C7D13933192AF6980619EE345E91EC2BB9D5CCA5C071E40926E54E8D
--R          DAE24196C0B2F8A0AAD60DBA57F5D4C8536262210C74F1
--R                                                   Type: HexadecimalExpansion
--E 4
--S 5 of 7
p := hex(1/4)*x**2 + hex(2/3)*x + hex(4/9)
--R
--R
--R            2     ____
--R   (5)  0.4x  + 0.Ax + 0.71C
--R                                       Type: Polynomial(HexadecimalExpansion)
--E 5
--S 6 of 7
q := D(p, x)
--R
--R
--R                 _
--R   (6)  0.8x + 0.A
--R                                       Type: Polynomial(HexadecimalExpansion)
--E 6
--S 7 of 7
g := gcd(p, q)
--R
--R
--R              _
--R   (7)  x + 1.5
--R                                       Type: Polynomial(HexadecimalExpansion)
--E 7
)spool
)lisp (bye)

———-

References^{[1] nothing}

5